EEVblog Electronics Community Forum

Products => Test Equipment => Topic started by: wlzed279 on August 23, 2023, 10:15:37 pm

Title: Tek 475 Bandwidth Limit Switch Not Functioning
Post by: wlzed279 on August 23, 2023, 10:15:37 pm
I have a newly acquired Tek 475 scope. The power rails are good and everything is functioning as it should with one exception. The bandwidth limit switch seem to be on the fritz. Left in the default position, I assume the 200mhz position, I get no trace at all. If I pull the switch to 100mhz, I get a somewhat noisy trace using a signal from the internal calibrator. I get the same results if i pull the switch to the 20mhz position. Interestingly, there is no change in the noise level between 100mhz and 20mhz.  I have cleaned the switch contacts using the paper/IPA method with no luck. Pushing the switch in I get a clean trace with no signal.

Not sure where to start on this. Should I be looking at U370/155-0091-00? How do I measure that IC to see if it is functioning properly.

Title: Re: Tek 475 Bandwidth Limit Switch Not Functioning
Post by: wasedadoc on August 24, 2023, 07:50:40 am
Why are you mentioning 20, 100 and 200 MHz positions of the switch? There is only one filter in that schematic. Pushing the switch changes from input signal to the trigger signal.

Do you have access to another 'scope?
Title: Re: Tek 475 Bandwidth Limit Switch Not Functioning
Post by: David Hess on August 24, 2023, 09:42:14 am
Something should sure change when the switch is used and it does not depend on U370.  The switch must still be dirty, or perhaps just broken.
Title: Re: Tek 475 Bandwidth Limit Switch Not Functioning
Post by: wlzed279 on August 24, 2023, 02:35:55 pm
Thanks David and wasedadoc for your replies. Attached is the corrected schematic. I do have another scope that I can use to trace the switch contacts. Looking at the schematic,  U370 supplies the full bandwidth signal and the switch directs the signal to the appropriate filters, correct? [attachimg=1]
Title: Re: Tek 475 Bandwidth Limit Switch Not Functioning
Post by: wasedadoc on August 24, 2023, 03:11:30 pm
Apply some known signal to the 475 scope input, set the coupling to AC and the Volt/div to the appropriate setting.

Check the 200 MHz position of the switch, as that is straight through and does not go though any filter components.  On the good scope you should see identical signals at A input and G output.  Similarly the signal at B input and H output should be identical.

Signals at A input and B input should be complementary (equal except inverted polarity.)  If not there is a problem in U370 or before that.

With the switch pushed and held in you should see, with the good scope, trigger signals at outputs G and H. If yes, but not on the 475 screen, there is a problem with the delay line or subsequent stages.  If no, and the outputs of U370 are not OK, problem is likely before the U370, ie at a point before the trigger signal is tapped off.
Title: Re: Tek 475 Bandwidth Limit Switch Not Functioning
Post by: David Hess on August 24, 2023, 03:18:57 pm
Looking at the schematic,  U370 supplies the full bandwidth signal and the switch directs the signal to the appropriate filters, correct?

That is correct.